Four resistors and two batteries are connected in a circuit, as shown.

1.522 2.5V WW w 112 MW 2.002 4.5V 4.002

a) Label the circuit and write down Kirchhoff’s rules!

b) Find the currents in the circuit.

c) Find the power delivered by the two batteries!

d) Find the power dissipated in each resistor!
Answer #1

12 Iz 4.5V - 2.5V 84.00 (a) sol:- at node A (b) IT 13 I + I₂ $2.00 loop 0 1.5 -2.5 + I + 2 I3 +1.51, ao 2-5 I +2 Iz = 2-5 lo(d) Power disse R, = 1.52 dissipated in resistors pa P= IR P = (1-043²x1.5 P = 1.6317 W Rr = 1 P = (1.043)*x/ P = 1.0878 ao =